# Simulation-based population inference of LISA’s Galactic binaries: Bypassing the global fit

### v1

@@ IV Where do constraints come from?
An important question regarding Galactic binaries is whether the information signature for population inference resides in the resolvable or the unresolvable sources. To shed light on this, we train and test two separate SBIs with identical network architectures on the two classes of sources. $\mathrm{SBI}_{\mathrm{Res}}$ ($\mathrm{SBI}_{\mathrm{Unres}}$) is trained to estimate the posteriors of the population from LISA’s channel-A data containing only instrumental noise and (un)resolvable DWDs. A DWD is deemed resolvable if its single detector optimal match-filter SNR $\rho_{\text{opt}}$ is greater than 7. $\rho_{\text{opt}}$ is computed assuming a fiducial noise PSD with $\mathcal{A}=\mathcal{O}=1$, augmented by the stochastic DWD background as parametrized in Ref. LDC2022 . The results presented below and the conclusions drawn do not qualitatively depend on this SNR threshold.
@@
Passing this data to both $\mathrm{SBI}_{\mathrm{Res}}$ and $\mathrm{SBI}_{\mathrm{Unres}}$, Fig. 9 compares their A channel posteriors. All parameters, barring high-frequency noise component $\mathcal{O}$, are better constrained by the resolvable subpopulation. Moreover, in comparison to $\mathrm{SBI}_{\mathrm{Unres}}$, the posteriors from $\mathrm{SBI}_{\mathrm{Res}}$ are more alike those from the total population in Fig. 6. The $\mathrm{SBI}_{\mathrm{Unres}}$ posteriors of the mass distribution parameters, Galactic scale radius, and separation index are far more diffuse in comparison to those of $\mathrm{SBI}_{\mathrm{Res}}$. These results, along with comparable $\log_{10}N,\,\mathcal{A}$ posteriors, suggest that the resolvable sources form a sufficiently representative subsample of the total population to produce informative posteriors. The improved posteriors of noise parameter $\mathcal{O}$ in $\mathrm{SBI}_{\mathrm{Unres}}$ as compared to $\mathrm{SBI}_{\mathrm{Res}}$ is likely due to a more accurate estimation of the noise characteristic at higher frequencies. This likely stems from the fact that in the case of the former, there are no sharp fluctuations caused by resolvable DWDs, especially at higher frequencies, which might otherwise be misattributed to noise. It is important to note that the conclusions drawn here are conditioned on the specific summary statistics used; alternatively, a different summary that may capture additional features from the unresolved sources could alter this interpretation.
@@ V Conclusion
In this work, we have introduced an amortized SBI framework to extract the population properties of Galactic DWDs from LISA frequency (or time) series data, bypassing the traditional and computationally expensive global-fit cataloging procedure.
We integrate a fast, GPU-accelerated forward simulator of the DWD population and its associated GW signals to generate several hundred thousand training examples. A conditional normalizing flow is then trained on summary statistics that capture both the collective signature of the unresolved sub-population and the features of the individually resolvable sources. The summarizer takes as input the absolute amplitude and real parts of the frequency spectrum, compressing the large number ($\mathcal{O}(10^{7})$) of frequency values into a much smaller set of summary statistics ($\mathcal{O}(6\times 10^{3})$). This enables rapid amortized posterior estimation of key population parameters, including the primary mass distribution, the Milky Way’s disk and bulge scale lengths, the bulge to disk fraction, the DWD separation distribution power-law index, and the total number of DWDs in the Galaxy.
Importantly, our analysis utilizes information from both the relatively few ($\mathcal{O}(10^{4})$) resolved sources and the vastly more numerous ($\mathcal{O}(10^{7})$) unresolved foreground sources. By using the full data, we naturally mitigate the selection biases that typically affect hierarchical inference based on finite detection catalogs, as in standard global-fit methods that focus only (at first) on resolvable sources. Moreover, once trained and applied to real LISA data, our SBI framework can provide instantaneous population-level constraints as new observations are collected, enabling continuously updated, real-time population inference throughout the mission.
Our results show that the calibrated SBI approach delivers rapid posterior estimates with (approximately) unbiased coverage, successfully recovering the injected population parameters with percent-level precision for the primary mass distribution, separation index, and total number of sources, and with tens-of-percent precision for the Galactic disk scales. Crucially, by including both amplitude and phase information in the summary statistics, we are able to break degeneracies that would otherwise leave the spatial-scale parameters prior-dominated.
@@
An important caveat of parametric reconstruction techniques—including SBI as well as tractable likelihood methods—is that the inferred posteriors are only as accurate as the forward model used to simulate the data; that is, the posterior probability is predicated on the assumption that the forward model is correct. It is therefore essential to perform inference using a variety of forward simulators to assess the robustness of the results. To this end, we emphasize that our framework is inherently flexible and can readily incorporate additional source classes (e.g., MBHBs, stBHBs, EMRIs) as well as different noise features, including non-Gaussianity and time-dependent artifacts, such as data gaps and glitches.
Future work may explore the use of SBI performed sequentially in time as LISA observations accumulate, providing real-time updates on source and noise parameter inference. Finally, in future iterations of the method, we may explore different or refined summary statistics, which can be learned via neural autoencoders to improve sensitivity to subtle population features.
@@ Acknowledgements.
We thank Chantal Pitte and Neil Cornish for their useful suggestions and comments, and Valeriya Korol for providing the code for the population synthesis. We acknowledge support from the European Union’s H2020 ERC Consolidator Grant “GRavity from Astrophysical to Microscopic Scales” (Grant No. GRAMS-815673, to E.B. and R.S.), the European Union’s Horizon ERC Synergy Grant “Making Sense of the Unexpected in the Gravitational-Wave Sky” (Grant No. GWSky-101167314, to E.B. and R.S.), the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing” (to E.B.), and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855 (to E.B.). R.T. acknowledges co-funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 Project FAIR “Future Artificial Intelligence Research”. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. R.T. is partially supported by the Fondazione ICSC, Spoke 3 “Astrophysics and Cosmos Observations”, Piano Nazionale di Ripresa e Resilienza Project ID CN00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19)” - Next Generation EU (NGEU). This work has been supported by the Agenzia Spaziale Italiana (ASI), Project n. 2024-36-HH.0, “Attività per la fase B2/C della missione LISA”. Computational work has been made possible through SISSA-CINECA and CINECA-INFN agreements providing access to resources on LEONARDO at CINECA.MB. N.K. acknowledges support from the CNES for the exploration of LISA science. N.K. gratefully acknowledges support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing and data-processing resources needed for this work.
@@ Appendix A Primary mass fit
Instead of the Gaussian mixture model in Maoz+2018 (requiring 8 free parameters), we model the primary mass $m_{1}$ of the DWD as a truncated Lorentzian (Cauchy) probability distribution (two free parameters) given by
@@
where $m_{\mathrm{0}}$ is the mass corresponding to the peak of the distribution and $m_{\mathrm{\gamma}}$ corresponds to the half-width half-maximum value. We sample $m_{\mathrm{0}}$ from a uniform distribution. For better representation of both narrow and flat mass distributions, we sample $m_{\mathrm{\gamma}}$ from a log-uniform prior ($\log_{10}m_{\mathrm{\gamma}}$ sampled uniformly). In Fig. 10, we compare our fit with that of the observation data and the Gaussian mixture model from Maoz+2018 , showing a reasonable overlap of all three distributions.
Figure 10: Comparison of the eight-parameter Gaussian mixture fit Kepler+2015 and our two-parameter Lorentzian fit of the observed white dwarf mass distribution.
@@
Reshape (Flatten) + Dropout - $(B,16384)$ Dropout probability $=0.2$
Fully Connected + ReLU + Dropout $(B,16384)$ $(B,256)$ Dropout probability $=0.2$
@@ Appendix C Network architecture
In this section, we describe the architecture of the SBI’s conditional normalizing flow network. The normalizing flow NF is conditioned on the context vector, a latent representation of the data extracted by a convolutional neural network CNN_Encoder.
@@ Appendix D Calibrated posteriors
Following the temperature scaling procedure outlined in §$\S$ II.4, we resample the uncalibrated posterior samples to obtain the calibrated distribution. Each uncalibrated sample $i$, with value $\bm{\Lambda}_{i}$ and associated probability $p_{\mathrm{S}BI,\,i}$, is assigned a weight $w_{i}=p_{\mathrm{S}BI,\,i}\,^{1-1/T^{*}}$. In Fig. 12, we compare the two distributions. The calibrated posteriors are generally narrower in comparison to the uncalibrated due to the underconfident coverage shown in the PP plot of Fig. 5.

### v2

@@ IV Where do constraints come from?
An important question regarding Galactic binaries is whether the information signature for population inference resides in the resolvable or the unresolvable sources. To shed light on this, we train and test two separate SBIs with identical network architectures on the two classes of sources. $\mathrm{SBI}_{\mathrm{Res}}$ ($\mathrm{SBI}_{\mathrm{Unres}}$) is trained to estimate the posteriors of the population from LISA's channel-A data containing only instrumental noise and (un)resolvable DWDs. A DWD is deemed resolvable if its single detector optimal match-filter SNR $\rho_{\text{opt}}$ is greater than 7. $\rho_{\text{opt}}$ is computed assuming a fiducial noise PSD with $\mathcal{A}=\mathcal{O}=1$, augmented by the stochastic DWD background as parametrized in Ref. LISA Data Challenge working group (2022). The results presented below and the conclusions drawn do not qualitatively depend on this SNR threshold.
@@
Passing this data to both $\mathrm{SBI}_{\mathrm{Res}}$ and $\mathrm{SBI}_{\mathrm{Unres}}$, Fig. 9 compares their A channel posteriors. All parameters, barring high-frequency noise component $\mathcal{O}$, are better constrained by the resolvable subpopulation. Moreover, in comparison to $\mathrm{SBI}_{\mathrm{Unres}}$, the posteriors from $\mathrm{SBI}_{\mathrm{Res}}$ are more alike those from the total population in Fig. 6. The $\mathrm{SBI}_{\mathrm{Unres}}$ posteriors of the mass distribution parameters, Galactic scale radius, and separation index are far more diffuse in comparison to those of $\mathrm{SBI}_{\mathrm{Res}}$. These results, along with comparable $\log_{10}N,\,\mathcal{A}$ posteriors, suggest that the resolvable sources form a sufficiently representative subsample of the total population to produce informative posteriors. The improved posteriors of noise parameter $\mathcal{O}$ in $\mathrm{SBI}_{\mathrm{Unres}}$ as compared to $\mathrm{SBI}_{\mathrm{Res}}$ is likely due to a more accurate estimation of the noise characteristic at higher frequencies. This likely stems from the fact that in the case of the former, there are no sharp fluctuations caused by resolvable DWDs, especially at higher frequencies, which might otherwise be misattributed to noise.
Finally we also note that Fig. 9 shows a slight negative correlation between the posteriors of $m_{\mathrm{0}}$ and $\alpha$, a negative correlation with $\log_{10}N$. These correlations are more prominent at lower SNRs and hence not seen in the joint population of Fig. 6. DWD populations from different mass distributions occupy different frequency regions (c.f. the frequency histogram in Fig. 2). Populations with a higher median mass (i.e., larger $m_{\mathrm{0}}$) have lower GW frequencies. Due to the lower PSD of LISA at these frequencies (c.f. the PSD in Fig. 2), this results in lower SNR. Thus, the correlation arises because of the similar effect that increasing $\log_{10}N$ and decreasing $m_{\mathrm{0}}$ has on the SNR of the sources.
It is important to note that the conclusions drawn here are conditioned on the specific summary statistics used; a different summary that captures additional features from the unresolved sources could alter this interpretation. Our compression method summarizes individually resolvable source peaks and the unresolvable power spectrum within a frequency bin by a linear fit, the standard deviation, and outlier power. Summaries with smaller frequency bins (greater resolution) and/or non-linear functions might present greater fidelity to the detector data. Inference from high-fidelity summaries, assuming a sufficiently expressive SBI and adequate training data, would be more accurate, potentially with better coverage before calibration. This is especially true in the analysis of GBs, given that a large fraction of the information appears to come from the resolvable peaks. However, with increasingly larger summaries, the number of network parameters increases rapidly, requiring proportionally large GPU VRAM. The degree of data compression is thus a tradeoff between fidelity towards input data and computational limitations.
@@ V Conclusion
In this work, we have introduced an amortized SBI framework to extract the population properties of Galactic DWDs from LISA frequency (or time) series data, without the traditional, computationally expensive DWD source-by-source extraction of the global-fit.
We integrate a fast, GPU-accelerated forward simulator of the DWD population and its associated GW signals to generate several hundred thousand training examples. A conditional normalizing flow is then trained on summary statistics that capture both the collective signature of the unresolved sub-population and the features of the individually resolvable sources. The summarizer takes as input the absolute amplitude and real parts of the frequency spectrum, compressing the large number ($\mathcal{O}(10^{7})$) of frequency values into a much smaller set of summary statistics ($\mathcal{O}(6\times 10^{3})$). This enables rapid amortized posterior estimation of key population parameters, including the primary mass distribution, the Milky Way's disk and bulge scale lengths, the bulge to disk fraction, the DWD separation distribution power-law index, and the total number of DWDs in the Galaxy.
Importantly, our analysis utilizes information from both the relatively few ($\mathcal{O}(10^{4})$) resolved sources and the vastly more numerous ($\mathcal{O}(10^{7})$) unresolved foreground sources. By using the full data, we naturally mitigate the selection biases that typically affect hierarchical inference based on finite detection catalogs, as in standard global-fit methods that focus only (at first) on resolvable sources. In the context of Galactic binary population inference, our proof of concept promises to complement the global-fit, at a fraction of the computation time. Moreover, once trained and applied to real LISA data, our SBI framework can provide instantaneous population-level constraints as new observations are collected, enabling continuously updated, real-time population inference throughout the mission.
Our results show that the calibrated SBI approach delivers rapid posterior estimates with (approximately) unbiased coverage, successfully recovering the injected population parameters with percent-level precision for the primary mass distribution, separation index, and total number of sources, and with tens-of-percent precision for the Galactic disk scales. Crucially, by including both amplitude and phase information in the summary statistics, we can break degeneracies that would otherwise leave the spatial-scale parameters prior-dominated.
@@
An important caveat of parametric reconstruction techniques—including SBI as well as tractable likelihood methods—is that the inferred posteriors are only as accurate as the forward model used to simulate the data; that is, the posterior probability is predicated on the assumption that the forward model is correct. It is therefore essential to perform inference using a variety of forward simulators to assess the robustness of the results. To this end, we emphasize that our framework is inherently flexible and can readily incorporate additional source classes (e.g., MBHBs, stBHBs, EMRIs) as well as un-modeled sources different noise features, including non-Gaussianity and time-dependent artifacts, such as data gaps and glitches. Modeling these features is necessary for both simulation-based and global fit methods.
Future work may explore the use of SBI performed sequentially in time as LISA observations accumulate, providing real-time updates on source and noise parameter inference. Finally, in future iterations of the method, we may explore different or refined summary statistics, which can be learned via neural autoencoders to improve sensitivity to subtle population features, or different SBI methods such as truncated marginal neural ratio estimation (see e.g. Refs. Bhardwaj et al. (2023); Alvey et al. (2024); Cole et al. (2025) for applications to GW data analysis).
@@ Acknowledgements.
We thank Chantal Pitte and Neil Cornish for their useful suggestions and comments, and Valeriya Korol for providing the code for the population synthesis. We acknowledge support from the European Union’s H2020 ERC Consolidator Grant ``GRavity from Astrophysical to Microscopic Scales'' (Grant No. GRAMS-815673, to E.B. and R.S.), the European Union’s Horizon ERC Synergy Grant ``Making Sense of the Unexpected in the Gravitational-Wave Sky'' (Grant No. GWSky-101167314, to E.B. and R.S.), the PRIN 2022 grant ``GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing'' (to E.B.), and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855 (to E.B.). R.T. acknowledges co-funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 Project FAIR ``Future Artificial Intelligence Research''. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. R.T. is partially supported by the Fondazione ICSC, Spoke 3 ``Astrophysics and Cosmos Observations'', Piano Nazionale di Ripresa e Resilienza Project ID CN00000013 ``Italian Research Center on High-Performance Computing, Big Data and Quantum Computing'' funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di ``campioni nazionali di R&S (M4C2-19)'' - Next Generation EU (NGEU). This work has been supported by the Agenzia Spaziale Italiana (ASI), Project n. 2024-36-HH.0, ``Attività per la fase B2/C della missione LISA''. Computational work has been made possible through SISSA-CINECA and CINECA-INFN agreements providing access to resources on LEONARDO at CINECA.MB. N.K. acknowledges support from the CNES for the exploration of LISA science. N.K. gratefully acknowledges support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing and data-processing resources needed for this work.
@@ Appendix A Primary mass fit
Instead of the Gaussian mixture model in Maoz et al. (2018) (requiring 8 free parameters), we model the primary mass $m_{1}$ of the DWD as a truncated Lorentzian (Cauchy) probability distribution (two free parameters) given by
@@
where $m_{\mathrm{0}}$ is the mass corresponding to the peak of the distribution and $m_{\mathrm{\gamma}}$ corresponds to the half-width half-maximum value. We sample $m_{\mathrm{0}}$ from a uniform distribution. For better representation of both narrow and flat mass distributions, we sample $m_{\mathrm{\gamma}}$ from a log-uniform prior ($\log_{10}m_{\mathrm{\gamma}}$ sampled uniformly). In Fig. 10, we compare our fit with that of the observation data and the Gaussian mixture model from Maoz et al. (2018), showing a reasonable overlap of all three distributions.
Figure 10: Comparison of the eight-parameter Gaussian mixture fit Kepler et al. (2015) and our two-parameter Lorentzian fit of the observed white dwarf mass distribution.
@@
Reshape (Flatten) + Dropout - $(B,16384)$ Dropout probability =0.2=0.2
Fully Connected + ReLU + Dropout $(B,16384)$ $(B,256)$ Dropout probability =0.2=0.2
@@ Appendix C Network architecture
In this section, we describe the architecture of the SBI's conditional normalizing flow network. The normalizing flow NF is conditioned on the context vector, a latent representation of the data extracted by a convolutional neural network CNN_Encoder.
@@ Appendix D Calibrated posteriors
Following the temperature scaling procedure outlined in $\S$ II.4, we resample the uncalibrated posterior samples to obtain the calibrated distribution. Each uncalibrated sample $i$, with value $\bm{\Lambda}_{i}$ and associated probability $p_{\mathrm{S}BI,\,i}$, is assigned a weight $w_{i}=p_{\mathrm{S}BI,\,i}\,^{1-1/T^{*}}$. In Fig. 12, we compare the two distributions. The calibrated posteriors are generally narrower in comparison to the uncalibrated due to the underconfident coverage shown in the PP plot of Fig. 5.